a)    English Class

 

For the host IP address, identify the prefix

192.168.10.1/

Identify the LSB (least significant bit)

 

Identify the Subnet ID to which the host belongs

 

Identify the broadcast address

 

Identify the usable range of host IP addresses

 


1
Expert's answer
2021-07-08T01:32:44-0400

The given IP 192.168.10.1/

If the number of hosts is 25 then the size of the block which fulfills the requirement "=2^5"

For the block size of 32, the number of host bits "=2^5" , hence host bits will be 5, which is also called least significant bits.

If hosts bit are 5 then number of network bit = 32-5 = 27 network bits

Number of network bits is 27, then sub-net mask = 8 | 8 | 8 | 3

"=255.255.255.224"

So, prefix 192.168.10.1/27

Least significant bit =5

The belonging sub-net = 192.168.10.0

Broadcast address = 192.168.10.31

Usable range of host IP = 192.168.10.1 to 192.168.10.30